# Underwater SONAR Image Classification and Analysis using LIME-based Explainable Artificial Intelligence

### v1

@@ Abstract
Index Terms:
Imaging SONAR, Underwater object classification, Transfer learning, Explainable AI (XAI), Local Interpretable Model-Agnostic Explanation (LIME).
I Introduction
@@
The forthcoming sections of this paper are organized as follows: The discussion of SONAR imaging and related work is presented in Section II. The development of the custom dataset used for training is described in Section III. Classification using pre-trained models and different Explainable Artificial Intelligence (XAI) methodologies, with a special focus on LIME and SP-LIME is described in Sections IV and V, respectively. The experimental findings are showcased in Section VI. Finally, discussion and conclusion are drawn in Section VII.
II Background and Related work
@@
II-A SONAR working Principle
@@
II-B Underwater Object Classification
@@
II-C Explainable AI for SONAR Image Analysis
@@ III Dataset for Learning
III-A Data Acquisition
@@
    The primary dataset, referred to as Seabed Objects KLSG  [24], comprises 1190 Side Scan SONAR images with a total of 385 images of shipwrecks, 36 images of drowning victims, 62 images of planes, 129 images of mines, and 578 images of the seafloor. The authors gathered these real SONAR images with the assistance of various commercial SONAR suppliers, including Lcocean [56], Klein Martin  [57], and EdgeTech [58] accumulated over a period of ten years. A few classes with a total of 1171 images with a variety of dimensions, along with augmented images excluding mines, are accessible to the public through the link [59], with the purpose of supporting academic research in this particular domain.
@@
TABLE I: Dataset samples: First row: class name; second row: number of images per class; the remaining corresponding rows are the sample images from that particular class.
By combining the aforementioned datasets, we curated our tailor-made dataset, which comprises a total of 2989 side scan SONAR images with 1334 fish images, 34 human images, 167 mine images, 123 plane images, 578 seafloor images, and 753 ship wreck images. Some sample images of our custom dataset, together with the image count for each label, are provided in Table I. All images in the dataset are obtained from the originally captured images without any preprocessing.
III-B Data Splitting
@@
IV Methodology: Transfer Learning-Based Image Classification, Data Balancing and LIME-based Explanation
@@
IV-A Transfer Learning
Transfer learning facilitates the rapid prototyping of new machine learning models by leveraging the weights and biases of pre-trained models from a source task [67]. This transfer of weights, essentially the transfer of knowledge, offers notable benefits [68] in several ways, not only in speeding up the convergence process during training but also in establishing a higher initial accuracy. This is especially advantageous when dealing with smaller datasets, demonstrating the efficiency and effectiveness of transfer learning [8].
@@ IV-A Transfer Learning
IV-B Handling Data Imbalance
As explained in Section III-B, there exists a class imbalance problem in the dataset due to high variance in the number of samples per class. This issue is tackled through a dual strategy: Oversampling and Undersampling techniques  [69] (refer Fig.3). Oversampling refers to the technique of adding more images to minority classes, whereas undersampling refers to the technique of randomly discarding some images from the majority class [70]. In our case, the majority classes are fishes, seafloor, and ships, while the minority classes are plane, human, and mine. To address the scarcity of plane, human, and mine images. We employ image augmentation techniques such as flipping, rotation, cropping, adjusting the brightness, contrast, random erasing, noise injection, mixing images, adding blur, sharpness, and colour jittering to generate synthetic images, as an instance of oversampling [71] (Refer to Fig.4). Conversely, to harmonise the abundance of ship, fish, and seafloor images, we randomly discard some with equal probabilities to balance image numbers for each class as an undersampling approach [72],  [73]. The number of images before and after augmentation is mentioned in Table II.
@@ IV-B Handling Data Imbalance
TABLE II: Number of Images/Class: First column: number of images before augmentation; & Second column: number of images after augmentation.
@@
IV-D Local interpretable Model-Agnostic Explanations (LIME) as an XAI tool
@@
$\pi_{x}(z)=\exp\left(-\frac{D(x,z)^{2}}{\sigma^{2}}\right)$ (2)
@@
IV-E Submodular Picks Local interpretable Model-Agnostic Explanations (SP-LIME) for images
The classifier may represent the image as a tensor with three colour channels per pixel, but an interpretable representation for our image classification task might be a binary vector indicating the “presence” or “absence” of a contiguous patch of similar pixels (a super-pixel). The process of explaining model predictions through LIME starts with the deliberate selection of an instance, or a group of instances (in our case, its super-pixels), that are worthy of interpretation. The selected instance’s features are then subtly altered to introduce perturbations, resulting in a diverse set of samples. Subsequently, the black-box model generates predictions for these altered samples.
@@ V-B Evaluation Protocols
VI Experimental Results
In this section, the experimental results of our study are presented. A detailed classification analysis of the different CNN benchmark models using different sampling strategies (random vs. stratified) and data balancing schemes are provided in Section VI-A and Section VI-B. Further, the explainer predictions of XAI models LIME and SP-LIME are detailed, along with various ablation studies and state-of-the-art comparisons in Sections  VI-CVI-D, and VI-E, respectively.
VI-A Classification Model-Random Sampling
As mentioned in Section III-B , we divide the dataset into three segments—train, validation, and test using random sampling approach in which images from the dataset are chosen at random. The forthcoming section explains the outcomes of the unbalanced data vs balanced data-based classification.
VI-A1 Unbalanced Data
The random sampling approach on an unbalanced dataset leads to an inconsistent number of images per class in all the splits, which is also non-linear (Refer to Table III). The customised deep learning models, initialised with pre-trained weights, undergo rigorous training and evaluation to assess their effectiveness in our classification task. Fig. 5 and Table IV summarises the accuracy achieved with its epoch count by each model during the training process. In particular, VGG16 -95.75%, VGG19 - 95.31%, ResNet50 - 92.18%, Inception V3 - 96.65%, DenseNet121 - 97.32%, MobileNetV2 - 96.42%, Xception - 96.20%, InceptionResNetV2 exhibited a high accuracy of 97.76%, DenseNet201 - 97.09%, NASNetLarge - 95.75% and NASNetMobile of 96.20% are obtained on the test dataset.
TABLE III: Image counts in the train, validation, and test sets utilizing random sampling approach with unbalanced data.
@@ VI-A1 Unbalanced Data
TABLE IV: Accuracy of Deep Learning Models developed utilizing random sampling approaches with unbalanced data.
@@
TABLE V: Classification Report for Unbalanced Model using random sampling approach-InceptionResNetV2
Class Precision Recall F1-Score Accuracy
Fish 1.00 1.00 1.00 1.00
Human 1.00 1.00 1.00 1.00
Mine 1.00 1.00 1.00 1.00
Plane 0.91 0.59 0.71 0.90
Seafloor 1.00 1.00 1.00 1.00
Ship 0.95 0.99 0.97 0.94
Although the accuracies are in an acceptable range, from the classification report (Refer to Table V and Fig. 7(a) it is observed that the test data is suboptimal for some classes such as plane, mine, and human. compared to majority classes of fish, shipwreck, and seafloor images. Such model testing with imbalanced data results in a biased model towards the majority class. To overcome this, data balancing technique i.e. using an equal number of images from each class to train the model to possibly balance the model, as mentioned in  [69] is envisaged.
VI-A2 Balanced Data
The predictive ability of the models developed in the previous case scenario (unbalanced data) is undoubtedly questionable due to the limited number of test images in the minority class and the abundance of test images in the majority class. To address this issue and achieve data balance across different classes, we employ oversampling and undersampling techniques, as extensively discussed in Section IV-B. As a result, an equal number of 498 images per class is achieved. Further, random sampling approach is utilized to split the dataset into train, validate, and test sets. Hence the number of images per class will be balanced across the splits, as depicted in Table VI. The process of image classification using transfer learning is performed again using this newly balanced data. The accuracy achieved by each model during the training process with balanced data along with its epoch count is summarised in Table VII and in Fig. 6. These results highlight the performance of DenseNet201, which achieved an accuracy of 98.88%. The other backbone models achieved VGG16 - 95.53%, VGG19 - 95.08%, ResNet50 - 83.92%, Inception V3 - 97.32%, MobileNetV2 - 96.65%, Xception - 97.54%, InceptionResNetV2 - 96.42%, DenseNet121 - 98.43%, NASNetLarge - 97.32% and NASNetMobile of 96.87% accuracies, respectively on the test dataset.
TABLE VI: Image counts in the train, validation, and test sets utilizing random sampling approach with balanced data.
@@ VI-A2 Balanced Data
TABLE VII: Accuracy of Deep Learning Models developed utilizing random sampling approaches with balanced data.
@@
TABLE VIII: Classification Report for Balanced Model using random sampling approach-DenseNet201.
Class Precision Recall F1-Score Accuracy
Fish 1.00 1.00 1.00 1.00
Human 0.97 0.96 0.97 0.98
Mine 1.00 1.00 1.00 1.00
Plane 0.97 0.95 0.96 0.97
Seafloor 0.97 1.00 0.99 0.97
Ship 0.95 0.97 0.96 0.94
Although there is a small drop in accuracy after balancing, the performance across all classes is consistent and well-balanced. Specifically, referring to Table V, the accuracy of the unbalanced model for the class labels ‘plane’ is 90%, Whereas, with the balanced model- DenseNet201 classification report, the accuracy has improved to 97%. However, some other class accuracies dropped slightly, e.g. human and seafloor (accuracies dropped from 100% to 98%). However, the model is generalized across all the classes (Refer Table VIII and Fig. 7(b)). This demonstrates a substantial improvement of balanced models on individual class performance compared to the performance of unbalanced models.
(a) InceptionResNetV2 - Unbalanced Model
(b) DenseNet201 - Balanced Model
VI-B Classification Model-Stratified Sampling
The accuracy of the models developed in the previous Section VI-A leveraged random sampling techniques. As a result, the data sample representation from each class varies in both training and evaluation. To address this issue, an alternative sampling strategy stratified sampling techniques is employed [85]. As mentioned in Section III-B, we divide the dataset into three segments— train, validation, and test using stratified sampling techniques, which ensure that images from all classes are selected with equal probability. The outcomes of the unbalanced data vs balanced data are discussed below:
VI-B1 Unbalanced Data
The stratified sampling approach on an unbalanced dataset leads to a consistent number of images per class in all the splits, which is also linear (Refer to Table IX).
TABLE IX: Image counts in the train, validation, and test sets utilizing stratified sampling approach with unbalanced data.
@@ VI-B1 Unbalanced Data
The customised deep learning models, initialised with pre-trained weights, undergo rigorous training and evaluation to assess their effectiveness in our classification task. As depicted in the Fig. 8, the following Table X summarises the accuracy achieved by each model, VGG16 achieved 96.65%, VGG19 - 96.65%, ResNet50 - 93.50%, Inception V3 - 97.10%, DenseNet121 - 98.44%, MobileNetV2, and DenseNet201 exhibited a accuracy of 97.55%, Xception - 97.32 %, InceptionResNetV2 - 97.99, NASNetMobile -97.32, and NASNetLarge - 96.21%.
TABLE X: Accuracy of Deep Learning Models developed utilizing stratified sampling approaches with unbalanced data.
@@
Although the accuracy is improved compared to the models developed using random sampling approach, the performance of individual classes are still lower compared to the balanced models developed using random sampling techniques (Refer Fig. 10(a)). By comparing the class-wise accuracies in Table V and Table XI, we can observe a significant increase in individual class performance. The unbalanced model from the random sampling approach achieved an accuracy of 90% for the class ”plane”, whereas the counterpart result of the same unbalanced data through a stratified sampling approach achieves an accuracy of 93%, which is considerably higher but not equivalent to the accuracy of other classes.
TABLE XI: Classification Report for unbalanced Model using stratified sampling approach-DenseNet121.
Class Precision Recall F1-Score Accuracy
Fish 1.00 1.00 1.00 1.00
Human 1.00 0.60 0.75 1.00
Mine 1.00 1.00 1.00 1.00
Plane 0.93 0.78 0.85 0.93
Seafloor 1.00 0.99 0.99 1.00
Ship 0.94 0.99 0.97 0.94
VI-B2 Balanced Data
The stratified approach improved the overall accuracy, but the individual class accuracy is still lower. Hence, the predictive ability of the models developed above utilizing stratified sampling approach is still undoubtedly questionable due to the limited number of test images in the minority class and the abundance of test images in the majority class. To address this issue and to achieve data balance across different classes, oversampling and undersampling techniques are carried out, as extensively discussed in Section IV-B. As a result, 498 images per class is achieved, and stratified sampling approach is utilized in spitting the dataset into train, validate, and test sets. The number of images per class after splitting get balanced, as mentioned in Table XII.
TABLE XII: Image counts in the train, validation, and test sets utilizing stratified sampling approach with balanced data.

### v2

@@ Abstract
1 Introduction
@@
The forthcoming sections of this paper are organized as follows: The discussion of SONAR imaging and related work is presented in Section 2. The development of the custom dataset used for training is described in Section 3. Classification using pre-trained models and different Explainable Artificial Intelligence (XAI) methodologies, with a special focus on LIME and SP-LIME is described in Sections 4 and 5, respectively. The experimental findings are showcased in Section 6. Finally, discussion and conclusion are drawn in Section 7.
2 Background and Related work
@@
2.1 SONAR working Principle
@@
2.2 Underwater Object Classification
@@
2.3 Explainable AI for SONAR Image Analysis
@@ 3 Dataset for Learning
Table 1: Dataset samples: First row: class name; second row: number of images per class; the remaining corresponding rows are the sample images from that particular class.
3.1 Data Acquisition
@@
    The primary dataset, referred to as Seabed Objects KLSG  [33], comprises 1190 Side Scan SONAR images with a total of 385 images of shipwrecks, 36 images of drowning victims, 62 images of planes, 129 images of mines, and 578 images of the seafloor. The authors gathered these real SONAR images with the assistance of various commercial SONAR suppliers, including Lcocean, Klein Martin, and EdgeTech accumulated over a period of ten years. A few classes with a total of 1171 images with a variety of dimensions, along with augmented images excluding mines, are accessible to the public through the link [34], with the purpose of supporting academic research in this particular domain.
@@
By combining the aforementioned datasets, we curated our tailor-made dataset, which comprises a total of 2989 side scan SONAR images with 1334 fish images, 34 human images, 167 mine images, 123 plane images, 578 seafloor images, and 753 ship wreck images. Some sample images of our custom dataset, together with the image count for each label, are provided in Table 1. All images in the dataset are obtained from the originally captured images without any preprocessing.
3.2 Data Splitting
@@
4 Methodology: Transfer Learning-Based Image Classification, Data Balancing and LIME-based Explanation
@@
4.1 Transfer Learning
Transfer learning facilitates the rapid prototyping of new machine learning models by leveraging the weights and biases of pre-trained models from a source task [14]. This transfer of weights, essentially the transfer of knowledge, offers notable benefits in several ways, not only in speeding up the convergence process during training but also in establishing a higher initial accuracy. This is especially advantageous when dealing with smaller datasets, demonstrating the efficiency and effectiveness of transfer learning [10].
@@ 4.1 Transfer Learning
4.2 Handling Data Imbalance
As explained in Section 3.2, there exists a class imbalance problem in the dataset due to high variance in the number of samples per class. This issue is tackled through a dual strategy: Oversampling and Undersampling techniques  [36] (refer Fig.3). Oversampling refers to the technique of adding more images to minority classes, whereas undersampling refers to the technique of randomly discarding some images from the majority class [46]. In our case, the majority classes are fishes, seafloor, and ships, while the minority classes are plane, human, and mine. To address the scarcity of plane, human, and mine images. We employ image augmentation techniques such as flipping, rotation, cropping, adjusting the brightness, contrast, random erasing, noise injection, mixing images, adding blur, sharpness, and colour jittering to generate synthetic images, as an instance of oversampling [69] (Refer to Fig.4). Conversely, to harmonise the abundance of ship, fish, and seafloor images, we randomly discard some with equal probabilities to balance image numbers for each class as an undersampling approach [44],  [37]. The number of images before and after augmentation is mentioned in Table 2.
@@ 4.2 Handling Data Imbalance
Table 2: Number of Images/Class: First column: number of images before augmentation; & Second column: number of images after augmentation.
@@
4.4 Local interpretable Model-Agnostic Explanations (LIME) as an XAI tool
@@
$\vspace{-0.15cm}\pi_{x}(z)=\exp\left(-\frac{D(x,z)^{2}}{\sigma^{2}}\right)$ (2)
@@
4.5 Submodular Picks Local interpretable Model-Agnostic Explanations (SP-LIME) for images
The classifier may represent the image as a tensor with three colour channels per pixel, but an interpretable representation for our image classification task might be a binary vector indicating the “presence" or “absence" of a contiguous patch of similar pixels (a super-pixel). The process of explaining model predictions through LIME starts with the deliberate selection of an instance, or a group of instances (in our case, its super-pixels), that are worthy of interpretation. The selected instance’s features are then subtly altered to introduce perturbations, resulting in a diverse set of samples. Subsequently, the black-box model generates predictions for these altered samples.
@@ 5.2 Evaluation Protocols
6 Experimental Results
In this section, the experimental results of our study are presented. A detailed classification analysis of the different CNN benchmark models using different sampling strategies (random vs. stratified) and data balancing schemes are provided in Section 6.1 and Section 6.2. Further, the explainer predictions of XAI models LIME and SP-LIME are detailed, along with various ablation studies and state-of-the-art comparisons in Sections  6.36.4, and 6.5, respectively.
6.1 Classification Model-Random Sampling
As mentioned in Section 3.2 , we divide the dataset into three segments—train, validation, and test using random sampling approach in which images from the dataset are chosen at random. The forthcoming section explains the outcomes of the unbalanced data vs balanced data-based classification.
6.1.1 Unbalanced Data
The random sampling approach on an unbalanced dataset leads to an inconsistent number of images per class in all the splits, which is also non-linear (Refer to Table 3). The customised deep learning models, initialised with pre-trained weights, undergo rigorous training and evaluation to assess their effectiveness in our classification task. Fig. 6 and Table 5 summarises the accuracy achieved with its epoch count by each model during the training process. In particular, VGG16 -95.75%, VGG19 - 95.31%, ResNet50 - 92.18%, Inception V3 - 96.65%, DenseNet121 - 97.32%, MobileNetV2 - 96.42%, Xception - 96.20%, InceptionResNetV2 exhibited a high accuracy of 97.76%, DenseNet201 - 97.09%, NASNetLarge - 95.75% and NASNetMobile of 96.20% are obtained on the test dataset.
Table 3: Image counts in the train, validation, and test sets utilizing random sampling approach with unbalanced data.
@@ 6.1.1 Unbalanced Data
Table 4: Classification Report for Unbalanced Model using random sampling approach-InceptionResNetV2
Class Precision Recall F1-Score Accuracy
Fish 1.00 1.00 1.00 1.00
Human 1.00 1.00 1.00 1.00
Mine 1.00 1.00 1.00 1.00
Plane 0.91 0.59 0.71 0.90
Seafloor 1.00 1.00 1.00 1.00
Ship 0.95 0.99 0.97 0.94
(a) InceptionResNetV2 - Unbalanced Model
(b) DenseNet201 - Balanced Model
Table 5: Accuracy of Deep Learning Models developed utilizing random sampling approaches with unbalanced data.
@@
Although the accuracies are in an acceptable range, from the classification report (Refer to Table 4 and Fig. 5(a) it is observed that the test data is suboptimal for some classes such as plane, mine, and human. compared to majority classes of fish, shipwreck, and seafloor images. Such model testing with imbalanced data results in a biased model towards the majority class. To overcome this, data balancing technique i.e. using an equal number of images from each class to train the model to possibly balance the model, as mentioned in  [36] is envisaged.
6.1.2 Balanced Data
The predictive ability of the models developed in the previous case scenario (unbalanced data) is undoubtedly questionable due to the limited number of test images in the minority class and the abundance of test images in the majority class. To address this issue and achieve data balance across different classes, we employ oversampling and undersampling techniques, as extensively discussed in Section 4.2. As a result, an equal number of 498 images per class is achieved. Further, random sampling approach is utilized to split the dataset into train, validate, and test sets. Hence the number of images per class will be balanced across the splits, as depicted in Table 6. The process of image classification using transfer learning is performed again using this newly balanced data. The accuracy achieved by each model during the training process with balanced data along with its epoch count is summarised in Table 8 and in Fig. 7. These results highlight the performance of DenseNet201, which achieved an accuracy of 98.88%. The other backbone models achieved VGG16 - 95.53%, VGG19 - 95.08%, ResNet50 - 83.92%, Inception V3 - 97.32%, MobileNetV2 - 96.65%, Xception - 97.54%, InceptionResNetV2 - 96.42%, DenseNet121 - 98.43%, NASNetLarge - 97.32% and NASNetMobile of 96.87% accuracies, respectively on the test dataset.
Table 6: Image counts in the train, validation, and test sets utilizing random sampling approach with balanced data.
@@ 6.1.2 Balanced Data
Table 7: Classification Report for Balanced Model using random sampling approach-DenseNet201.
Class Precision Recall F1-Score Accuracy
Fish 1.00 1.00 1.00 1.00
Human 0.97 0.96 0.97 0.98
Mine 1.00 1.00 1.00 1.00
Plane 0.97 0.95 0.96 0.97
Seafloor 0.97 1.00 0.99 0.97
Ship 0.95 0.97 0.96 0.94
Table 8: Accuracy of Deep Learning Models developed utilizing random sampling approaches with balanced data.
@@
Although there is a small drop in accuracy after balancing, the performance across all classes is consistent and well-balanced. Specifically, referring to Table 4, the accuracy of the unbalanced model for the class labels ‘plane’ is 90%, Whereas, with the balanced model- DenseNet201 classification report, the accuracy has improved to 97%. However, some other class accuracies dropped slightly, e.g. human and seafloor (accuracies dropped from 100% to 98%). However, the model is generalized across all the classes (Refer Table 7 and Fig. 5(b)). This demonstrates a substantial improvement of balanced models on individual class performance compared to the performance of unbalanced models.
6.2 Classification Model-Stratified Sampling
The accuracy of the models developed in the previous Section 6.1 leveraged random sampling techniques. As a result, the data sample representation from each class varies in both training and evaluation. To address this issue, an alternative sampling strategy stratified sampling techniques is employed [58]. As mentioned in Section 3.2, we divide the dataset into three segments— train, validation, and test using stratified sampling techniques, which ensure that images from all classes are selected with equal probability. The outcomes of the unbalanced data vs balanced data are discussed below:
6.2.1 Unbalanced Data
The stratified sampling approach on an unbalanced dataset leads to a consistent number of images per class in all the splits, which is also linear (Refer to Table 9).
Table 9: Image counts in the train, validation, and test sets utilizing stratified sampling approach with unbalanced data.
@@ 6.2.1 Unbalanced Data
Table 10: Classification Report for unbalanced Model using stratified sampling approach-DenseNet121.
Class Precision Recall F1-Score Accuracy
Fish 1.00 1.00 1.00 1.00
Human 1.00 0.60 0.75 1.00
Mine 1.00 1.00 1.00 1.00
Plane 0.93 0.78 0.85 0.93
Seafloor 1.00 0.99 0.99 1.00
Ship 0.94 0.99 0.97 0.94
The customised deep learning models, initialised with pre-trained weights, undergo rigorous training and evaluation to assess their effectiveness in our classification task. As depicted in the Fig. 8, the following Table 11 summarises the accuracy achieved by each model, VGG16 achieved 96.65%, VGG19 - 96.65%, ResNet50 - 93.50%, Inception V3 - 97.10%, DenseNet121 - 98.44%, MobileNetV2, and DenseNet201 exhibited a accuracy of 97.55%, Xception - 97.32 %, InceptionResNetV2 - 97.99, NASNetMobile -97.32, and NASNetLarge - 96.21%.
Table 11: Accuracy of Deep Learning Models developed utilizing stratified sampling approaches with unbalanced data.
@@
(a) DenseNet121 - Unbalanced Model
(b) DenseNet121 - Balanced Model
Although the accuracy is improved compared to the models developed using random sampling approach, the performance of individual classes are still lower compared to the balanced models developed using random sampling techniques (Refer Fig. 9(a)). By comparing the class-wise accuracies in Table 4 and Table 10, we can observe a significant increase in individual class performance. The unbalanced model from the random sampling approach achieved an accuracy of 90% for the class "plane", whereas the counterpart result of the same unbalanced data through a stratified sampling approach achieves an accuracy of 93%, which is considerably higher but not equivalent to the accuracy of other classes.
6.2.2 Balanced Data
The stratified approach improved the overall accuracy, but the individual class accuracy is still lower. Hence, the predictive ability of the models developed above utilizing stratified sampling approach is still undoubtedly questionable due to the limited number of test images in the minority class and the abundance of test images in the majority class. To address this issue and to achieve data balance across different classes, oversampling and undersampling techniques are carried out, as extensively discussed in Section 4.2. As a result, 498 images per class is achieved, and stratified sampling approach is utilized in spitting the dataset into train, validate, and test sets. The number of images per class after splitting get balanced, as mentioned in Table 12.
Table 12: Image counts in the train, validation, and test sets utilizing stratified sampling approach with balanced data.